# Hyperfocal subalgebras of hyperfocal abelian Frobenius blocks

### v1

@@ 5 A stable equivalence of Morita type
We still use the notation in the last two sections. Throughout this section, we assume that the block is a hyperfocal abelian Frobenius block. The main result in this section is the construction of a stable equivalence of Morita type between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$ under this assumption. To accomplish this, we need to glue the indecomposable endopermutation $k(D/Q)$-modules $V_{Q}$ for any $1\neq Q\leq D$ appearing in Proposition 4.11 together. The main technique is the result due to Puig [23, Proposition 3.6] (see also [17, Theorem 7.8.2]). For convenience, we present this result here.
@@
([23, Proposition 3.6]) Let $Z$ be a finite abelian $p$-group and $E$ a subgroup of $\mathrm{Aut}(Z)$. Let $\mathcal{C}$ be a nonempty upwardly closed $E$-stable set of subgroups of $Z$, which means that if $Q\leq R\leq Z$ with $Q$ in $\mathcal{C}$, then $R$ and $y(Q)$ belong to $\mathcal{C}$ fro all $y\in E$. For any subgroup $Q$ in $\mathcal{C}$ let $V_{Q}$ be an indecomposable endopermutation $k(Z/Q)$-module with vertex $Z/Q$. Suppose that if $Q,R$ belong to $\mathcal{C}$ such that $Q\leq R$, then $\mathrm{Defres}_{Z/R}^{Z/Q}(V_{Q})\cong V_{R}$, where $Z/R$ is identified with $(Z/Q)/(R/Q)$. Then there is an indecomposable endopermutation $kZ$-module $V$ with vertex $Z$ such that $\mathrm{Defres}_{Z/Q}^{Z}(V)\cong V_{Q}$ for any subgroup $Q$ in $\mathcal{C}$. In addition, if ${{}_{y}}(V_{Q})\cong V_{y(Q)}$ as $k(Z/y(Q))$-modules for all $y\in E$ and all $Q$ in $\mathcal{C}$, then ${{}_{y}}V\cong V$ as $kZ$-modules for all $y\in E$.
@@
In our setting, we can identify $N_{G}(D_{\delta})/C_{G}(D)$ with a subgroup of $\mathrm{Aut}(D)$. Denote $N_{G}(D_{\delta})/C_{G}(D)$ by $\tilde{N}$ and the image of $g\in N_{G}(D_{\delta})$ in $\tilde{N}$ by $\tilde{g}$. Let $\mathcal{C}$ be the set of subgroups of $D$ consisting of all nontrivial subgroups of $D$. Obviously, $\mathcal{C}$ is a nonempty upwardly $\tilde{N}-$stable set. Take two subgroups $Q$ and $R$ of $D$ with $Q\leq R$. We denote $\mathrm{Defres}_{D/R}^{D/Q}(V_{Q})$ by $W_{Q,R}$. Clearly, $W_{Q,R}$ is up to isomorphism a unique endopermutation $k(D/R)$-module with vertex $D/R$ such that $\bar{S}_{Q}(R)$ is isomorphic to $\mathrm{End}_{k}(W_{Q,R})$ as $D/R$-algebras. Here, $\bar{S}_{Q}$ is $\mathrm{End}_{k}(V_{Q})$.
@@
By Proposition 4.11, we have $\mathbb{A}_{\delta}(Q)\cong\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ and $\mathbb{A}_{\delta}(R)\cong\bar{S}_{R}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras. Since $D$ is abelian and $Q\leq R$, we have $\mathbb{A}_{\delta}(R)\cong(\mathbb{A}_{\delta}(Q))(R)$ as $D$-interior algebras. It is clear that $(\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD)(R)\cong\bar{S}_{Q}(R)\mathop{\otimes}\limits_{k}(kD)(R)\cong\bar{S}_{Q}(R)\mathop{\otimes}\limits_{k}kD$. Then we have $\bar{S}_{Q}(R)\mathop{\otimes}\limits_{k}kD\cong\bar{S}_{R}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras. By the uniqueness of $V_{R}$, we have $W_{Q,R}\cong V_{R}$ as $k(D/R)$-modules since $\bar{S}_{Q}(R)\cong\mathrm{End}_{k}(W_{Q,R})$.
Take an element $\tilde{g}$ of $\tilde{N}$. Clearly, $\mathrm{End}_{k}(_{g}V_{Q})\cong{{}_{g}}\bar{S}_{Q}$. Here, ${{}_{g}}\bar{S}_{Q}$ is the $D/{{}^{g}}Q$-interior algebra with being equal to $\bar{S}_{Q}$ as $k$-algebra and the group homomorphism from $D/{{}^{g}}Q$ to $({{}_{g}}\bar{S}_{Q})^{*}$ being just the composition of the group homomorphism from $D/{{}^{g}}Q$ to $D/Q$ induced by the conjugation of $g$ and the structural homomorphism from $D/Q$ to $\bar{S}_{Q}^{*}$. We use the similar notation for any $D$-interior algebra. Since $\mathbb{A}_{\delta}(Q)\cong\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras, we have
@@
as $D$-interior algebras. For $\tilde{g}\in\tilde{N}$, it can be written as $ux$ for some $u\in P$ and some $x\in N_{G}(D_{\delta})$ with $\tilde{x}\in E_{\mathfrak{h}}$. Borrowing the notation from the paragraph 3.1 and Proposition 4.12, we have $a_{u,\delta}\in A_{\delta}^{*}\cap j\mathbb{A}uj$ and $\mathfrak{e}_{x}\in N_{\mathbb{A}_{\delta}}(D)$ such that $a_{u,\delta}(wj)a_{u,\delta}^{-1}=(uwu^{-1})j$ and $\mathfrak{e}_{x}(wj)\mathfrak{e}_{x}^{-1}=(xwx^{-1})j$ for any $w\in D$. Denote $a_{u,\delta}\mathfrak{e}_{x}$ by $a_{ux,\delta}$. Then $a_{g,\delta}$ is in $N_{A_{\delta}}(D)\cap j\mathbb{A}uj$ and $a_{g,\delta}(wj)a_{g,\delta}^{-1}=(gwg^{-1})j$ for any $w\in D$. In particular, the conjugation action of $a_{g,\delta}^{-1}$ on $\mathbb{A}_{\delta}$ induces a $D$-interior algebra isomorphism $\mathbb{A}_{\delta}\cong{{}_{g}}\mathbb{A}_{\delta}$. Therefore,
@@
By Lemmas 5.1 and 5.3, we can get an indecomposable endopermutation $kD$-module $\bar{V}$ with vertex $D$ such that $(\mathrm{End}_{k}(\bar{V}))(Q)\cong\bar{S}_{Q}$ as $D$-interior algebras for any nontrivial subgroup $Q$ of $D$ and ${{}_{g}}\bar{V}\cong\bar{V}$ as $kD$-modules for any $\tilde{g}\in\tilde{N}$. Furthermore, by [30, Theorem 14.2], there is an indecomposable endopermutation $\mbox{$\mathcal{O}$}D$-module $V$ with vertex $D$ such that $k\mathop{\otimes}\limits_{\mathcal{O}}V\cong\bar{V}$. Setting $S=\mathrm{End}_{\mathcal{O}}(V)$, then $S(Q)\cong\bar{S}_{Q}$ as $D$-interior algebras for any nontrivial subgroup $Q$ of $D$. In the following, we will show that $V$ can be chose to be $\tilde{N}-$stable, namely, ${{}_{g}}V\cong V$ as $kD$-modules for any $\tilde{g}\in\tilde{N}$.
@@
Take an element $\tilde{g}\in\tilde{N}$. Denote the structural homomorphism from $D$ to $S^{*}$ by $\phi$. Clearly, $\mathrm{End}_{\mathcal{O}}({{}_{g}}V)\cong{{}_{g}}S$ and the the structural homomorphism from $D$ to $({{}_{g}}S)^{*}$, denoted by ${{}_{g}}\phi$, mapping $w\in D$ to $\phi(gwg^{-1})$. Since $k\mathop{\otimes}\limits_{\mathcal{O}}{{}_{g}}V\cong{{}_{g}}\bar{V}$, we get $k\mathop{\otimes}\limits_{\mathcal{O}}{{}_{g}}V\cong k\mathop{\otimes}\limits_{\mathcal{O}}V$ as $kD$-modules. By [17, Proposition 7.3.12], $S$ is isomorphic to ${{}_{g}}S$ as $D$-algebras. Denote this isomorphism by $\psi$. Obviously, $\psi$ is an $\mathcal{O}$-automorphism of $S$. By the Skolem-Noether Theorem, $\psi$ is in fact an inner automorphism. In particular, $\psi$ preserves the determinants. Set $\mathrm{rank}_{\mathcal{O}}(V)=n$. By [17, Proposition 7.3.10 (i)], $n$ is coprime to $p$. Then the statements in this lemma follows from [29, Proposition (21.5)]. ∎
@@
Now we fix an indecomposable endopermutation $\mbox{$\mathcal{O}$}D$-module $V$ with vertex $D$ such that setting $S=\mathrm{End}_{\mathcal{O}}(V)$, $S(Q)\cong\bar{S}_{Q}$ for any nontrivial subgroup $Q$ of $D$ and $\mathrm{det}(w\cdot 1_{S})=1$ for any $w\in D$. Then by Lemma 5.5, we have ${{}_{g}}V\cong V$ as $\mbox{$\mathcal{O}$}D$-modules for any $\tilde{g}\in\tilde{N}$. In particular, $V$ is $E_{\mathfrak{h}}$-stable. By [17, Corollary 7.8.4], $V$ can be extended to an indecomposable $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-module $U$ such that $U$ has a vertex $D$ and a source $V$ and $\mathrm{Res}_{D}^{D\rtimes E_{\mathfrak{h}}}(U)\cong V$.
@@
We still denote $W$ by $V$ whenever this does not introduce any confusion. Now $S=\mathrm{End}_{\mathcal{O}}(V)$ is a $P$-stable primitive $L$-interior algebra. This means that for any $u\in P$ there exists an element $s_{u}$ in $S$ such that $s_{u}(x\cdot\mathrm{v})=(uxu^{-1})\cdot s_{u}(\mathrm{v})$ for any $x\in L$ and any $\mathrm{v}\in V$. In particular $s_{u}$ belongs to $S^{L}$ when $u$ is in $C_{P}(D)$. Since taking the inverse gives a group isomorphism from $S^{*}$ to $(S^{\circ})^{*}$, the opposite algebra $S^{\circ}$ of $S$ is also a $P$-stable primitive $L$-algebra.
@@
Recall from Proposition 4.12 that $\mbox{$\mathcal{O}$}L$ is a subalgebra of $\mathbb{A}_{\delta}$. So we can consider the $L$-interior algebra $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$. The following result gives the $\mbox{$\mathcal{O}$}(D\times D)$-module structure of $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$, which is similar to the one of $\mathbb{A}_{\delta}$ in Proposition 4.12.
@@
where $\mathbb{X}$ is isomorphic to a direct sum of modules of the form $\mathrm{Ind}_{\Delta_{\tilde{h}}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$ for some proper subgroup $Q$ of $D$ and some $\tilde{h}\in E_{\mathfrak{h}}$. In particular, $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ has a finite $D\times D$-stable $\mathcal{O}$-basis such that $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ is projective as a left $\mbox{$\mathcal{O}$}D$-module and as a right $\mbox{$\mathcal{O}$}D$-module.
@@
By Proposition 4.12, we have
$S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}\cong(\mathop{\bigoplus}\limits_{\tilde{g}\in E_{\mathfrak{h}}}\mathrm{Ind}_{\Delta_{\tilde{g}}(D)}^{D\times D}(\mathrm{Res}_{\Delta_{\tilde{g}}(D)}^{D\times D}(S^{\circ})))\oplus(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{X}),$
where $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{X}$ is isomorphic to a direct sum of modules of the form $\mathrm{Ind}_{\Delta_{\tilde{h}}(Q)}^{D\times D}(\mathrm{Res}_{\Delta_{\tilde{h}}(Q)}^{D\times D}(S^{\circ}))$ for some proper subgroup $Q$ of $D$ and some $\tilde{h}\in E_{\mathfrak{h}}$. For any $\tilde{g}\in E_{\mathfrak{h}}$, we denote by $s^{\circ}_{\tilde{g}}$ its image in $(S^{\circ})^{*}$. Let $\Omega^{\circ}$ be a $D$-stable $\mathcal{O}$-basis of $S^{\circ}$ containing $1_{S^{\circ}}$ as the unique $D$-invariant element. Then $\Omega^{\circ}\cdot(s^{\circ}_{\tilde{g}})^{-1}$ is a $\Delta_{\tilde{g}}(D)$-stable $\mathcal{O}$-basis of $S^{\circ}$ containing $(s^{\circ}_{\tilde{g}})^{-1}$ as the unique $\Delta_{\tilde{g}}(D)$-invariant element. We are done. ∎
@@
Let $e$ be a primitive idempotent of $(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})^{L}$ with $\mathrm{Br}_{D}(e)\neq 0$. Set $\mathbb{A}^{\prime}=e(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})e$. So $\mathbb{A}^{\prime}$ is a primitive $L$-interior algebra. In fact, $e$ is still primitive in $(\mathbb{A}^{\prime})^{D}$. Indeed, by [16, Corollary 5.5.9], every point of $D$ on $\mathbb{A}^{\prime}$ is local. Since $\mathbb{A}_{\delta}$ is a primitive $D$-interior algebra with $\mathbb{A}_{\delta}(D)\neq 0$, by [17, Theorem 7.4.1 (i)], there is a unique local point of $D$ on $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ with multiplicity $1$. Therefore, $e$ remains primitive in $(\mathbb{A}^{\prime})^{D}$. In the following, we collect some properties of $\mathbb{A}^{\prime}$.
@@
(ii) The $(\mathbb{A}^{\prime},\mathbb{A}_{\delta})$-bimodule $e\cdot(V^{*}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})$ induces a Morita equivalence between $\mathbb{A}^{\prime}$ and $\mathbb{A}_{\delta}$. Here, $V^{*}=\mathrm{Hom}(V,\mbox{$\mathcal{O}$})$ is the $\mathcal{O}$-dual module of $V$.

### v2

@@ 5 A stable equivalence of Morita type
We still use the notation in the last two sections. Throughout this section, we assume that the block is a hyperfocal abelian Frobenius block. The main result in this section is the construction of a stable equivalence of Morita type between $\mathbb{A}_{\delta}$ and $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$ under this assumption. To accomplish this, we need to glue the indecomposable endopermutation $k(D/Q)$-modules $V_{Q}$ for any $1\neq Q\leq D$ appearing in Proposition 4.13 together. The main technique is the result due to Puig [23, Proposition 3.6] (see also [16, Theorem 7.8.2]). For convenience, we present this result here.
@@
([23, Proposition 3.6]) Let $Z$ be a finite abelian $p$-group and $E$ a subgroup of $\mathrm{Aut}(Z)$. Let $\mathcal{C}$ be a nonempty upwardly closed $E$-stable set of subgroups of $Z$, which means that if $Q\leq R\leq Z$ with $Q$ in $\mathcal{C}$, then $R$ and $y(Q)$ belong to $\mathcal{C}$ for all $y\in E$. For any subgroup $Q$ in $\mathcal{C}$, let $V_{Q}$ be an indecomposable endopermutation $k(Z/Q)$-module with vertex $Z/Q$. Suppose that if $Q,R$ belong to $\mathcal{C}$ such that $Q\leq R$, then $\mathrm{Defres}_{Z/R}^{Z/Q}(V_{Q})\cong V_{R}$, where $Z/R$ is identified with $(Z/Q)/(R/Q)$. Then there is an indecomposable endopermutation $kZ$-module $V$ with vertex $Z$ such that $\mathrm{Defres}_{Z/Q}^{Z}(V)\cong V_{Q}$ for any subgroup $Q$ in $\mathcal{C}$. In addition, if ${{}_{y}}(V_{Q})\cong V_{y(Q)}$ as $k(Z/y(Q))$-modules for all $y\in E$ and all $Q$ in $\mathcal{C}$, then ${{}_{y}}V\cong V$ as $kZ$-modules for all $y\in E$.
@@
In our setting, we can identify $N_{G}(D_{\delta})/C_{G}(D)$, denoted by $\tilde{N}$, with a subgroup of $\mathrm{Aut}(D)$. For any $g\in N_{G}(D_{\delta})$, we denote its image in $\tilde{N}$ by $\tilde{g}$. Let $\mathcal{C}$ be the set of subgroups of $D$ consisting of all nontrivial subgroups of $D$. Obviously, $\mathcal{C}$ is a nonempty upwardly $\tilde{N}-$stable set. Take two subgroups $Q$ and $R$ of $D$ with $Q\leq R$. We denote $\mathrm{Defres}_{D/R}^{D/Q}(V_{Q})$ by $W_{Q,R}$. By definition, $W_{Q,R}$ is up to isomorphism the unique indecomposable endopermutation $k(D/R)$-module with vertex $D/R$ such that $\bar{S}_{Q}(R)$ is isomorphic to $\mathrm{End}_{k}(W_{Q,R})$ as $D/R$-algebras. Here, $\bar{S}_{Q}$ is $\mathrm{End}_{k}(V_{Q})$.
@@
By Proposition 4.13, we have $\mathbb{A}_{\delta}(Q)\cong\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ and $\mathbb{A}_{\delta}(R)\cong\bar{S}_{R}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras. Since $D$ is abelian and $Q\leq R$, we have $\mathbb{A}_{\delta}(R)\cong(\mathbb{A}_{\delta}(Q))(R)$ as $D$-interior algebras. It is clear that $(\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD)(R)\cong\bar{S}_{Q}(R)\mathop{\otimes}\limits_{k}(kD)(R)\cong\bar{S}_{Q}(R)\mathop{\otimes}\limits_{k}kD$. Then we have $\bar{S}_{Q}(R)\mathop{\otimes}\limits_{k}kD\cong\bar{S}_{R}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras. By the uniqueness of $V_{R}$ shown in Proposition 4.13, we have $W_{Q,R}\cong V_{R}$ as $k(D/R)$-modules since $\bar{S}_{Q}(R)\cong\mathrm{End}_{k}(W_{Q,R})$.
Take an element $\tilde{g}$ of $\tilde{N}$. Clearly, $\mathrm{End}_{k}(_{g}V_{Q})\cong{{}_{g}}\bar{S}_{Q}$. Here, ${{}_{g}}\bar{S}_{Q}$ is the $D/{{}^{g}}Q$-interior algebra with being equal to $\bar{S}_{Q}$ as $k$-algebras and the group homomorphism from $D/{{}^{g}}Q$ to $({{}_{g}}\bar{S}_{Q})^{\times}$ being just the composition of the group homomorphism from $D/{{}^{g}}Q$ to $D/Q$ induced by the conjugation action of $g$ and the structural homomorphism from $D/Q$ to $\bar{S}_{Q}^{\times}$. We use the similar notation for any $D$-interior algebra. Since $\mathbb{A}_{\delta}(Q)\cong\bar{S}_{Q}\mathop{\otimes}\limits_{k}kD$ as $D$-interior algebras, we have
@@
as $D$-interior algebras. For $\tilde{g}\in\tilde{N}$, it can be written as $\tilde{u}\tilde{x}$ for some $u\in P$ and some $x\in N_{G}(D_{\delta})$ with $\tilde{x}\in E_{\mathfrak{h}}$. Borrowing the notation from the paragraph 3.1 and Proposition 4.16, there are elements $a_{u,\delta}\in A_{\delta}^{\times}\cap j\mathbb{A}uj$ and $\mathfrak{e}_{x}\in N_{\mathbb{A}_{\delta}}(D)$ such that $a_{u,\delta}(wj)a_{u,\delta}^{-1}=(uwu^{-1})j$ and $\mathfrak{e}_{x}(wj)\mathfrak{e}_{x}^{-1}=(xwx^{-1})j$ for any $w\in D$. Set $a_{g,\delta}=a_{u,\delta}\mathfrak{e}_{x}$. Then $a_{g,\delta}$ is in $N_{A_{\delta}}(D)\cap j\mathbb{A}uj$ and $a_{g,\delta}(wj)a_{g,\delta}^{-1}=(gwg^{-1})j$ for any $w\in D$. In particular, the conjugation action of $a_{g,\delta}^{-1}$ on $\mathbb{A}_{\delta}$ induces a $D$-interior algebra isomorphism $\mathbb{A}_{\delta}\cong{{}_{g}}\mathbb{A}_{\delta}$. Therefore,
@@
By Lemmas 5.1 and 5.3, we can get an indecomposable endopermutation $kD$-module $\bar{V}$ with vertex $D$ such that $(\mathrm{End}_{k}(\bar{V}))(Q)\cong\bar{S}_{Q}$ as $D$-interior algebras for any nontrivial subgroup $Q$ of $D$ and ${{}_{g}}\bar{V}\cong\bar{V}$ as $kD$-modules for any $\tilde{g}\in\tilde{N}$. Furthermore, by [30, Theorem 14.2], there is an indecomposable endopermutation $\mbox{$\mathcal{O}$}D$-module $V$ with vertex $D$ such that $k\mathop{\otimes}\limits_{\mathcal{O}}V\cong\bar{V}$. Setting $S=\mathrm{End}_{\mathcal{O}}(V)$, then $S(Q)\cong\bar{S}_{Q}$ as $D$-interior algebras for any nontrivial subgroup $Q$ of $D$. In the following, we will show that $V$ can be chose to be $\tilde{N}-$stable, namely, ${{}_{g}}V\cong V$ as $\mbox{$\mathcal{O}$}D$-modules for any $\tilde{g}\in\tilde{N}$.
@@
Take an element $\tilde{g}$ in $\tilde{N}$. Denote the structural homomorphism from $D$ to $S^{\times}$ by $\phi$. Clearly, $\mathrm{End}_{\mathcal{O}}({{}_{g}}V)\cong{{}_{g}}S$ and the the structural homomorphism from $D$ to $({{}_{g}}S)^{\times}$, denoted by ${{}_{g}}\phi$, mapping $w\in D$ to $\phi(gwg^{-1})$. Since $k\mathop{\otimes}\limits_{\mathcal{O}}{{}_{g}}V\cong{{}_{g}}\bar{V}$, we get $k\mathop{\otimes}\limits_{\mathcal{O}}{{}_{g}}V\cong k\mathop{\otimes}\limits_{\mathcal{O}}V$ as $kD$-modules. By [16, Proposition 7.3.12], $S$ is isomorphic to ${{}_{g}}S$ as $D$-algebras. Denote this isomorphism by $\psi$. Obviously, $\psi$ is an $\mathcal{O}$-automorphism of $S$. By the Skolem-Noether Theorem, $\psi$ is in fact an inner automorphism. In particular, $\psi$ preserves the determinants. Set $\mathrm{rank}_{\mathcal{O}}(V)=n$. By [16, Proposition 7.3.10 (i)], $n$ is coprime to $p$. Then the statements in this lemma follows from [29, Proposition (21.5)]. ∎
@@
Now we fix an indecomposable endopermutation $\mbox{$\mathcal{O}$}D$-module $V$ with vertex $D$ such that setting $S=\mathrm{End}_{\mathcal{O}}(V)$, $S(Q)\cong\bar{S}_{Q}$ for any nontrivial subgroup $Q$ of $D$ and $\mathrm{det}(w\cdot 1_{S})=1$ for any $w\in D$. Then by Lemma 5.5, we have ${{}_{g}}V\cong V$ as $\mbox{$\mathcal{O}$}D$-modules for any $\tilde{g}\in\tilde{N}$. In particular, $V$ is $E_{\mathfrak{h}}$-stable. By [16, Corollary 7.8.4], $V$ can be extended to an indecomposable $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{h}})$-module $U$ such that $U$ has a vertex $D$ and a source $V$ and $\mathrm{Res}_{D}^{D\rtimes E_{\mathfrak{h}}}(U)\cong V$.
@@
We still denote $W$ by $V$ whenever this does not introduce any confusion. Now $S=\mathrm{End}_{\mathcal{O}}(V)$ is a $P$-stable primitive $L$-interior algebra. This means that for any $u\in P$ there exists an element $s_{u}$ in $S^{\times}$ such that $s_{u}(x\cdot\mathrm{v})=(uxu^{-1})\cdot s_{u}(\mathrm{v})$ for any $x\in L$ and any $\mathrm{v}\in V$. In particular $s_{u}$ belongs to $S^{L}$ when $u$ is in $C_{P}(D)$. Since taking the inverse gives a group isomorphism from $S^{\times}$ to $(S^{\circ})^{\times}$, the opposite algebra $S^{\circ}$ of $S$ is also a $P$-stable primitive $L$-algebra.
@@
Recall from Proposition 4.16 that $\mbox{$\mathcal{O}$}L$ can be viewed as a subalgebra of $\mathbb{A}_{\delta}$. So we can consider the $L$-interior algebra $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$. The following result gives the $\mbox{$\mathcal{O}$}(D\times D)$-module structure of $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$, which is similar to the one of $\mathbb{A}_{\delta}$ in Proposition 4.16.
@@
where $\mathbb{X}$ is isomorphic to a direct sum of modules of the form $\mathrm{Ind}_{\Delta_{\tilde{h}}(Q)}^{D\times D}(\mbox{$\mathcal{O}$})$ for some proper subgroup $Q$ of $D$ and some $\tilde{h}\in E_{\mathfrak{h}}$. Here, we denote the subgroup $\{(u,g^{-1}ug)\mid u\in Q\}$ of $D\times D$ by $\Delta_{\tilde{g}}(Q)$ for any $\tilde{g}\in E_{\mathfrak{h}}$ and any subgroup $Q$ of $D$. In particular, $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ has a finite $D\times D$-stable $\mathcal{O}$-basis such that $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ is projective as a left $\mbox{$\mathcal{O}$}D$-module and as a right $\mbox{$\mathcal{O}$}D$-module.
@@
By Proposition 4.16, as $\mbox{$\mathcal{O}$}(D\times D)$-modules,
$S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}\cong(\mathop{\bigoplus}\limits_{\tilde{g}\in E_{\mathfrak{h}}}\mathrm{Ind}_{\Delta_{\tilde{g}}(D)}^{D\times D}(\mathrm{Res}_{\Delta_{\tilde{g}}(D)}^{D\times D}(S^{\circ})))\oplus(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}X),$
where $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}X$ is isomorphic to a direct sum of indecomposable modules of the form $\mathrm{Ind}_{\Delta_{\tilde{h}}(Q)}^{D\times D}(\mathrm{Res}_{\Delta_{\tilde{h}}(Q)}^{D\times D}(S^{\circ}))$ for some proper subgroup $Q$ of $D$ and some $\tilde{h}\in E_{\mathfrak{h}}$. For any $\tilde{g}\in E_{\mathfrak{h}}$, we denote by $s^{\circ}_{\tilde{g}}$ its image in $(S^{\circ})^{\times}$. Let $\Omega^{\circ}$ be a $D$-stable $\mathcal{O}$-basis of $S^{\circ}$ containing $1_{S^{\circ}}$ as the unique $D$-invariant element. Then $\Omega^{\circ}\cdot(s^{\circ}_{\tilde{g}})^{-1}$ is a $\Delta_{\tilde{g}}(D)$-stable $\mathcal{O}$-basis of $S^{\circ}$ containing $(s^{\circ}_{\tilde{g}})^{-1}$ as the unique $\Delta_{\tilde{g}}(D)$-invariant element. Then the statements in this lemma follow from the existence of this $\mathcal{O}$-basis of $S^{\circ}$ and the decomposition above. ∎
@@
Let $e$ be a primitive idempotent of $(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})^{L}$ with $\mathrm{Br}_{D}(e)\neq 0$. Set $\mathbb{A}^{\prime}=e(S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})e$. So $\mathbb{A}^{\prime}$ is a primitive $L$-interior algebra. In fact, $e$ is still primitive in $(\mathbb{A}^{\prime})^{D}$. Indeed, by [15, Corollary 5.5.9], every point of $D$ on $\mathbb{A}^{\prime}$ is local. Since $\mathbb{A}_{\delta}$ is a primitive $D$-interior algebra with $\mathbb{A}_{\delta}(D)\neq 0$, by [16, Theorem 7.4.1 (i)], there is a unique local point of $D$ on $S^{\circ}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta}$ with multiplicity $1$. Therefore, $e$ remains primitive in $(\mathbb{A}^{\prime})^{D}$. In the following, we collect some properties of the $L$-interior algebra $\mathbb{A}^{\prime}$.
@@
(ii) The $\mathbb{A}^{\prime}$-$\mathbb{A}_{\delta}$-bimodule $e\cdot(V^{*}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})$ induces a Morita equivalence between $\mathbb{A}^{\prime}$ and $\mathbb{A}_{\delta}$. Here, the right $\mathbb{A}_{\delta}$-module structure of $e\cdot(V^{*}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}_{\delta})$ is given by the right multiplication in $\mathbb{A}_{\delta}$.